### v1

# Balancing Diversity and Risk in LLM Sampling: How to Select Your Method and Parameter for Open-Ended Text Generation

Abstract
Sampling-based decoding strategies have been widely adopted for Large Language Models (LLMs) in numerous applications, which target a balance between diversity and quality via temperature tuning and tail truncation (e.g., top-k and top-p sampling). Considering the high dynamic range of the candidate next-tokens given different prefixes, recent studies propose to adaptively truncate the tail of LLM’s predicted distribution. Although improved results haven been reported with these methods on open-ended text generation tasks, the results are highly dependent on the curated truncation parameters and exemplar text. In this paper, we propose a systematic way to estimate the intrinsic capacity of a truncation sampling method by considering the trade-off between diversity and risk at each decoding step, based on our collected prefix tree which preserves the context of a full sentence. Our work provides a comprehensive comparison between existing truncation sampling methods, as well as their recommended parameters as a guideline for users.
Balancing Diversity and Risk in LLM Sampling: How to Select Your Method and Parameter for Open-Ended Text Generation
Yuxuan Zhou1,3, Margret Keuper1,2, Mario Fritz3 1University of Mannheim, Germany 2Max Planck Institute for Informatics, Saarland Informatics Campus, Germany 3CISPA Helmholz Center for Information Security, Germany zhouyuxuanyx@gmail.com, keuper@mpi-inf.mpg.de, fritz@cispa.de
1 Introduction
Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023); Team et al. (2023) have demonstrated incredible performance in a variety of applications, and the reliability of decoding strategies has become a critical concern, especially where diverse and coherent samples are desired. Previous works have revealed that likelihood-maximization such as beam search Fan et al. (2018); Holtzman et al. (2020); Welleck et al. (2020); Meister et al. (2022) produces degenerated text which contains repetitive loops and incoherent context, particularly in the open-ended tasks. Therefore, sampling-based decoding strategies, e.g., Top-p Holtzman et al. (2020) and Top-k sampling Radford et al. (2018); Fan et al. (2018), have been widely adopted. The balance between diversity and quality of the generated text could be adjusted by tuning the temperature and truncation position to some extend, but requires non-trivial trial and error.
Recent studies Basu et al. (2021); Zhu et al. (2024); Hewitt et al. (2022); Meister et al. (2023) proposed adaptive tail truncation mechanisms based on different criteria or assumptions, which maintain an allowed set of tokens with a flexible size according to the given prefix. To validate the effectiveness of a sampling method, they are often compared through extrinsic evaluation based on open-ended text generation applications. For example, story generation Fan et al. (2018) and document continuation Merity et al. (2017). Various metrics Welleck et al. (2020); Meister et al. (2023); Pillutla et al. (2021); Gao et al. (2021) have been adopted to consider different aspects of the generated text.
We reveal that there exist two underlying issues in the current evaluation setup, which might hinder the assessment of a sampling method’s practical significance in real-world applications:
The improvement of one method against another may be simply due to a better tuned parameter for the targeted exemplar text: the performance of sampling methods is sensitive to their parameters, and parameter sweep is often operated on a extremely sparse grid due to the high computation cost. This is especially problematic considering the non-linear dependency between performance and parameters.
Users are agnostic to the optimal parameters in real-world applications: Practically speaking, users often pick parameters based on their own need for the compromise between diversity and quality, after few number of tryouts. There exits no universal optimal hyperparamters in different scenarios and users are agnostic to the optimal hyerparameters for their own tasks.
The above issues exactly indicate the need for an evaluation that allows for estimating the theoretical capacity of a sampling method, independent of hyperparamter tuning. Moreover, the second issue additionally highlights the need for identifying the sweet spots of existing sampling methods, which could serve as a general guideline for parameter selection when applying a method.
Based on the above analysis, we propose a systematic way to assess the inherent adaptability of a sampling method in different contexts. First, we rearrange Wikipedia-English 111https://dumps.wikimedia.org/ data in the form of a word-level prefix tree structure, or the so-called Trie Fredkin (1960); Ghasemi et al. (2019). As shown in fig. 3, all possible words that appear after a given prefix in the dataset are collected together as the child nodes, and their preceding word is regarded as the parent node. Starting from "Begin of Sequence" and collecting the child nodes recursively, we are able to transform the full dataset into a single prefix tree. It is noteworthy that a n-gram Trie Jurafsky (2000) tends to produce overestimated data support size given a prefix Bengio et al. (2000), due to the lost of contextual information outside the truncation window, as shown in fig. 1. In a similar spirit to Ding et al. (2024), we intentionally construct the prefix tree with only sentence-starting n-grams to preserve the context of a full sentence. thus we refer to the collected data as Context-Preserving Trie (CP-Trie).
Since the child nodes of the ancestral prefix define the data support, the recall of a sampling method w.r.t. data support is able to be computed. Though the training datasets of modern LLMs are significantly larger than Wikipedia dataset, such a data support serve as a reasonable lower-bound, especially when only sentence-level context are considered for evaluation, see section 4.1 for more details.
Given the context-preserving prefix tree, we are able to evaluate the theoretical capacity of a sampling method, by examining the amount of tokens within and out of the data support with varying truncation parameter values. As can be seen in fig. 2, the optimal truncation position vary drastically given different prefixes and Top-k sampling could be regarded as a baseline method with zero adaptability. Therefore, an effective adaptive truncation mechanism is supposed to better follow such a variation, so that improved diversity can be achieved without harming the quality.
In summary, the contributions of our paper are as follows:
We establish an intrinsic evaluation benchmark based on the Context-Preserving Prefix tree data, which allows for estimating the theoretical capacity of different sampling decoding methods via our proposed diversity and stability metrics.
We conduct a comprehensive comparison of existing sampling approaches on our proposed benchmark, which serve as a guideline for choosing the truncation sampling methods and their corresponding parameter selection in real-world applications.
2 Related Work
In this section, we summarize the recently proposed sampling-based decoding strategies and the widely adopted benchmarks as well as metrics for open-ended text generation.
2.1 Sampling-based Decoding Methods
Vanilla sampling suffers from the risk of obtaining incoherent tokens, thus truncation of the tail distribution has been heavily discussed to alleviate such an issue, e.g., Top-k (Radford et al. (2018); Fan et al. (2018)) and Top-p sampling (Holtzman et al. (2020)). However, a fixed k or p is problematic when considering the high dynamic range of next reasonable tokens, as pointed out in more recent studies on adaptive sampling methods: Mirostat Basu et al. (2021) is proposed based on Zipf statistics and the assumption of a steady perplexity during generation. Hewitt et al. (2022) introduce -sampling which dismisses the tokens with low probabilities in the tail of the predicted distribution based on absolute and relative thresholds. Locally Typical Sampling Meister et al. (2023) assumes the generated text to retain a similar entropy rate to that of human-generated text. Adaptive Decoding Zhu et al. (2024) proposes to keep the entropy of the truncated distribution close to the original entropy. Although these approaches have been empirically evaluated on open-ended text generation tasks with curated truncation parameters and exemplar text, there still lacks a comprehensive comparison of their adaptability in more general cases.
2.2 Evaluation of Sampling-based Decoding
Benchmarks The commonly adopted benchmarks include story generation with WritingPrompts dataset Fan et al. (2018), document continuation with WikiText-103 dataset Merity et al. (2017) and abstractive summarization on the CNN/DAILYMAIL dataset Nallapati et al. (2016). These benchmarks suffer from the problem of limited exemplar text, which fails to capture the diverse nature of human language.
Statistical metrics are mostly based on n-gram statistics and focus on a single aspect, such as Repetition Welleck et al. (2020), Diversity Meister et al. (2023), Semantic coherence Gao et al. (2021), Zipf’s coefficient Holtzman et al. (2020) (simple Unigram rank-frequency statistics) and Self-BLEU Zhu et al. (2018).
Exemplar-based metrics dominate the evaluation of sampling-based decoding methods. As observed by Fan et al. (2018); Holtzman et al. (2020), lower perplexity of the generated text doesn’t necessarily indicate better quality. And Holtzman et al. (2020) suggested that the perplexity of the generated text is supposed to be close to that of the human text. MAUVE Pillutla et al. (2021) takes the trade-off between precision and recall into account, by comparing the learnt distribution from a text generation model to the distribution of human-written text using divergence frontiers. A recent study Shi et al. (2024a) provides a comprehensive evaluation on a large collection of tasks, which are mostly based on exemplar-based metrics.
3 Revisiting Truncation Sampling
3.1 Problem Formulation
Definition 3.1.
where denotes the allowed set comprising candidate next-tokens for a given prefix, and is the renormalization term.
Given the Context-Preserving Trie of a reference dataset, we could compute the estimate of the optimal allowed set as follows :
Definition 3.2.
Let be the allowed set after truncation given the prefix . The optimal allowed set corresponds to the allowed set with the minimum size, while covering the full data support. It is the solution to the following objective function:
3.2 Remaining Issues
We reveal three major issues in the evaluation of truncation sampling. We first summarize the problem of directly using probability as quality metric, then show the choice of truncation parameter has a significant impact on the evaluation. Finally, we reveal that a minor difference of Recall and Risk values may result in significant changes in diversity and quality of the generated text.
Unreliable Probability The probabilities of both the predicted and empirical distribution are not reliable for reflecting the quality of a text.
Word frequencies are average statistics across various topics, and assuming the optimal probabilities or the optimal ranking of each reasonable next token is ill-posed.
Impact of Parameter Selection We highlight the complexity and biases in parameter selection: Top-k and Top-p have constant upper bounds, i.e., the vocabulary size and 1, respectively. In contrast, the upper bounds of -sampling and adaptive sampling are dependent on LLM’s predicted distribution, because they truncate the tail distribution based on the likelihood of tokens and the slope of Min-Max scaled entropy, respectively. The importance of identifying the effective ranges of such parameters is also reflected in the authors’ choice of numeral digit for their proposed parameters. For example, is set to in Zhu et al. (2024) and is chosen from , and etc in Hewitt et al. (2022). In comparison, the adopted values for top-p sampling are merely two digits after zero, such as . Our analysis also shows the significance of identifying the sweet spots of different sampling methods.
The Butterfly Effect Although the top few samples possess the most probability mass, we reveal that a minor change in the size of the allowed set at each decoding step could lead to major differences in the quality of the generated text, due to LLMs’ auto-regressive nature. For example, the overall probability of at least one bad token to appear at the position in a sequence of length increases rapidly as increases (due to the product operator):
Similarly, if only probability mass is assigned to additional tokens at each step, for a sequence of length , there will be chances of obtaining extra diverse samples.
4 Method
In this section, we derive our method for evaluating different sampling-based decoding strategies. To circumvent the reliability issue of the probabilities we merely check whether the predicted next-token is in or out of the data support.
4.1 Probability-Independent Metrics
To quantify the diversity and quality of a sampling method based on CP-Trie, we define the Recall and Risk of a sampling method regarding a given prefix below:
Definition 4.1.
is dependent on the parameter selection for truncation, e.g., k value in top-k sampling. When the allowed set is smaller than the optimal one after truncation, Recall is smaller than one and Risk is regarded as zero. With further increased size of the allowed set, Recall reaches one but Risk emerges. Since the sizes of reasonable sets vary drastically for different prefixes, it is not possible to always retain the optimal allowed set with a predefined parameter. In this case, we reveal that the adaptability w.r.t. the varying size of data support of a sampling method indeed determines its effectiveness in real-world application.
Note that we ignore the risk of obtaining bad samples within the optimal allowed set, because such type of risk is unsolvable by truncation and is rather determined by the inherent capacity of the trained LLMs. However, such risk is less severe comparing to that introduced by inappropriate truncation, since LLMs exhibit a significant capability in predicting the next token Touvron et al. (2023); Achiam et al. (2023); Jiang et al. (2023); Team et al. (2023) and most bad samples reside in the tail distribution.
More importantly, our evaluation doesn’t rely on the empirical probability, which is biased and inaccurate due to limited dataset size or context window size. However, the tokens which appear in the dataset could be confidently regarded as reasonable, regardless of their actual probabilities. In addition, considering that temperature could change the flatness of distribution arbitrarily, we use ratio of token counts instead of probability mass to make the evaluation independent on temperature tuning and exemplar text. For a more detailed discussion with supporting examples, please refer to section A.2.
4.2 Parameter-Independent Evaluation
To eliminate the huge impact of parameter tuning on fair evaluation, we define the final diversity metric Recall at a given Risk level as follows:
Definition 4.2.
Analogously, a family of critical values such as can be easily defined.
From the definition, it can be seen that such a metric is no longer dependent on the selection of parameter, thus it reflects the genuine capacity of a sampling method regardless of parameter tuning. This allows for a fair comparison between different sampling methods, especially considering their drastically different effective ranges, as mentioned in section 1 and section 3.2.
4.3 The Priority of Low Variance in Risk
Despite that more diverse text is desired for many generation tasks, the minimum requirement of the text to be coherent and reasonable is mostly prioritized in real-world applications. Besides the average recall at a given risk level, it is noteworthy that a stable adaptive truncation mechanism is also preferred , i.e., the variance of risks should be kept as low as possible at the given average risk level.
Conjecture 4.3.
At a given average risk level, the total amount of risk when generating a sequence of length is reduced with decreased variance of the risks at each decoding step.
The above conjecture can be mainly understood by the auto-regressive generation process of LLMs, as shown in eq. 3. The in-distribution probability dependent on the risk at each decoding step and is minimized when the product of the probabilities is maximized at a given average risk level. We could infer that the sum of the in-distribution probabilities is approximated determined at a given average risk level. For simplification, we assume their sum is unchanged. According to AM-GM inequality, the maximum of the product is achieved when each individual component is equal to each other. Roughly speaking, the less variance of the probability masses at each step, the larger their product and thus the smaller the total risk is.
5 Experiment
In this section, we conduct evaluation of existing sampling-based decoding approaches on our re-collected EnWiki CP-Trie dataset. We aim to estimate the inherent adaptability of exiting sampling-based methods and the results could be used as references for the application of LLMs in open-ended tasks.
5.1 Data Collection
As mentioned above, the construction of the CP-Trie is a re-collection of an existing dataset. We apply the described procedure to the English subset of Wikipedia dataset and name the resulting dataset EnWiki CP-Trie. Although the core idea is straight-forward to understand, we elaborate the main design choices in the following:
Basic Unit There are many possible units for splitting the datasets into individual fragments, such as article, paragraph, sentence and n-grams. Constructing a tree based on articles or paragraphs may require a larger amount of data than the training data of LLMs to guarantee an adequate number of branches (because LLMs lean to interpolate), whereas the construction based on n-grams suffers from poor contextual information and are heavily biased towards common tuplets of n tokens regardless of the context. Therefore, we adopt sentence as the basic unit for our dataset, which guarantees a coherent context at sentence-level and requires a much smaller amount of data than the training data.
Filtering To avoid invalid words or rare proper names which are unreasonable for the model to predict, we exclude the sentences containing such words by checking their presence in the WORD LIST dataset, which is available on the website 222word-list dataset homepage. It contains a total amount of 354986 words and explicitly excludes proper names and compound words. Section titles are also excluded, because they are often incomplete sentences with poor contextual information.
Statistics Wikipedia-English dataset has a total number of articles, which results in EnWiki CP-Trie with number of leaves after conversion, as shown in fig. 4.
Storage The extracted prefix tree is implemented as a nested dictionary and saved in a single JSON file. Since each lookup at any depth has constant complexity, the retrieval from our dataset is highly efficient. Moreover, the dictionary is easily extendable if extra data are needed for a more accurate estimation of the full data support.
5.2 Evaluation Setup
Baselines Our evaluation includes Top-k sampling Radford et al. (2018); Fan et al. (2018), Top-p sampling Holtzman et al. (2020), -sampling Hewitt et al. (2022), Adaptive sampling Zhu et al. (2024) and Mirostat Basu et al. (2021) into comparison.
Evaluation Data To guarantee a tight lower bound of the ideal data support given different prefixes, we first sort the sub-nodes according to their total number of leaves at each depth, then we select the top 10 sub-trees with different sentence starting tokens for evaluation. Moreover, we keep the top 2 child nodes at each depth till depth 6, since the empirical data support becomes less adequate at large depth. This results in an evaluation set of prefixes with varying lengths in total.
Evaluation Metrics As discussed in section 4.2, we measure the improvement in diversity via the increase of average recall at a given risk level, and the reduction of the total risk in the auto-regressive process via the decrease of standard deviation at a given risk level, also referred to as stability.
LLMs To ensure that the obtained conclusion generalizes to different models, we adopt Llama-2 Touvron et al. (2023), Llama-3 Dubey et al. (2024), Mistral Jiang et al. (2023, 2024) families of different sizes and GPT-2-XL Radford et al. (2019) for comparison.
Implementation Our implementation mainly relies on Pytorch Paszke et al. (2017), HuggingFace Wolf et al. (2020) and OpenAI API 333https://pypi.org/project/openai/ library. We implement a truncation sampling method ourselves if the official implementation is not available. For all truncation methods, the minimum size of the allowed set is set to 1 to prevent breaking the sampling process.
| Model | Method | Avg. risk level 1 | Avg. risk level 5 | Avg. risk level 15 | ||||||
| Parameter | Risk Std Error | Recall | Parameter | Risk Std Error | Recall | Parameter | Risk Std Error | Recall | ||
| GPT2-XL | Top-k | 15 | 0.006 | 0.220 | 64 | 0.613 | 0.290 | 184 | 1.781 | 0.340 |
| Top-p | 0.5705 | 0.015 | 0.170 | 0.746 | 2.129 | 0.240 | 0.8555 | 6.210 | 0.338 | |
| Adaptive | 9.5e-4 | 0.006 | 0.252 | 1.1e-4 | 0.679 | 0.339 | 2.5e-05 | 2.241 | 0.413 | |
| Eta | 0.318 | 0.013 | 0.198 | 0.011 | 1.484 | 0.301 | 0.001 | 4.261 | 0.404 | |
| Mirostat | 4.425 | 0.005 | 0.236 | 5.9475 | 0.717 | 0.326 | 6.76 | 2.501 | 0.401 | |
| Llama-2-7b | Top-k | 14 | 0.126 | 0.226 | 61 | 0.587 | 0.296 | 177 | 1.722 | 0.369 |
| Top-p | 0.54 | 0.529 | 0.156 | 0.7665 | 2.331 | 0.254 | 0.9 | 6.208 | 0.400 | |
| Adaptive | 1.1e-3 | 0.154 | 0.257 | 1.4e-4 | 0.856 | 0.364 | 3.1e-5 | 2.966 | 0.470 | |
| Eta | 0.512 | 0.563 | 0.192 | 0.023 | 2.599 | 0.297 | 0.002 | 6.531 | 0.407 | |
| Mirostat | 4.253 | 0.133 | 0.236 | 5.82 | 0.650 | 0.349 | 6.628 | 2.286 | 0.474 | |
| Llama-2-70b | Top-k | 14 | 0.128 | 0.232 | 60 | 0.583 | 0.307 | 174 | 1.712 | 0.375 |
| Top-p | 0.6535 | 0.475 | 0.189 | 0.8465 | 2.136 | 0.316 | 0.9395 | 5.522 | 0.468 | |
| Adaptive | 0.0011 | 0.142 | 0.269 | 1.2e-4 | 0.796 | 0.374 | 2.3e-5 | 2.697 | 0.485 | |
| Eta | 0.092 | 0.304 | 0.236 | 0.003 | 1.590 | 0.378 | 2.1e-4 | 4.243 | 0.510 | |
| Mirostat | 4.16 | 0.135 | 0.238 | 5.7875 | 0.684 | 0.353 | 6.67 | 2.125 | 0.478 | |
| Llama-3-8B | Top-k | 14 | 0.128 | 0.228 | 59 | 0.576 | 0.290 | 172 | 1.701 | 0.346 |
| Top-p | 0.5395 | 0.451 | 0.154 | 0.736 | 2.061 | 0.224 | 0.855 | 5.770 | 0.326 | |
| Adaptive | 1.1e-3 | 0.167 | 0.260 | 1.7e-4 | 0.787 | 0.343 | 3.7e-5 | 2.685 | 0.418 | |
| Eta | 0.673 | 0.445 | 0.181 | 0.029 | 2.112 | 0.271 | 0.002 | 6.009 | 0.373 | |
| Mirostat | 4.24 | 0.139 | 0.230 | 5.8175 | 0.804 | 0.318 | 6.693 | 2.630 | 0.393 | |
| Llama-3-70B | Top-k | 14 | 0.127 | 0.230 | 60 | 0.581 | 0.295 | 173 | 1.695 | 0.352 |
| Top-p | 0.5695 | 0.502 | 0.158 | 0.758 | 2.386 | 0.237 | 0.8705 | 6.685 | 0.332 | |
| Adaptive | 1.1e-3 | 0.137 | 0.263 | 1.4e-4 | 0.787 | 0.353 | 3.16e-5 | 2.778 | 0.424 | |
| Eta | 0.37 | 0.137 | 0.263 | 0.014 | 2.231 | 0.295 | 0.001 | 6.265 | 0.398 | |
| Mirostat | 4.21 | 0.138 | 0.230 | 5.91 | 0.708 | 0.332 | 6.84 | 2.193 | 0.417 | |
| Mixtral-7B | Top-k | 14 | 0.126 | 0.224 | 62 | 0.596 | 0.297 | 181 | 1.759 | 0.364 |
| Top-p | 0.6565 | 0.539 | 0.194 | 0.8375 | 2.476 | 0.303 | 0.9315 | 6.315 | 0.447 | |
| Adaptive | 0.00105 | 0.152 | 0.260 | 1.2e-4 | 0.809 | 0.364 | 2.2e-5 | 2.757 | 0.466 | |
| Eta | 0.075 | 0.307 | 0.243 | 0.003 | 1.542 | 0.368 | 1.96e-4 | 4.712 | 0.505 | |
| Mirostat | 4.1825 | 0.141 | 0.236 | 5.8125 | 0.721 | 0.345 | 6.71 | 2.213 | 0.468 | |
| Mixtral-8x7B | Top-k | 15 | 0.134 | 0.229 | 63 | 0.598 | 0.301 | 183 | 1.757 | 0.366 |
| Top-p | 0.6505 | 0.535 | 0.192 | 0.8375 | 2.423 | 0.303 | 0.9325 | 6.139 | 0.456 | |
| Adaptive | 0.00105 | 0.148 | 0.265 | 1.1e-4 | 0.798 | 0.372 | 2.1e-5 | 2.802 | 0.476 | |
| Eta | 0.087 | 0.335 | 0.241 | 0.003 | 1.822 | 0.375 | 2.15e-4 | 4.922 | 0.506 | |
| Mirostat | 4.2775 | 0.143 | 0.238 | 5.845 | 0.710 | 0.346 | 6.6875 | 2.213 | 0.461 | |
5.3 Comparison at Different Risk Levels
In this section, we conduct a comprehensive study of different truncation sampling methods at different risk levels. As discussed in section 4.2, this allows for a fair comparison which is independent of parameter tuning. Moreover, we provide the corresponding parameters for each truncation sampling method at different risk levels, which could serve as user reference for the parameter selection of the compared methods.
As can be seen in table 1, different truncation sampling methods are compared at the average risk level of 1, 5, and 15 respectively. As discussed in section 4.1, our defined risk and recall metrics explicitly exclude the source of risk induced by a LLM’s capacity by design, thus similar parameter values correspond to the same risk level for most sampling methods across various model types and sizes. This exactly showcases the advantage of our evaluation being parameter-independent and sustainable to the rapid update of LLMs. Among the evaluated methods, Eta-sampling Hewitt et al. (2022) is the most sensitive to the changes of model type and size especially at risk levels of 1 and 5, which might hinder its practical significance at a low risk level.
Regarding diversity, i.e., the average recall at the same average risk level, Adaptive sampling Zhu et al. (2024) and Mirostat Basu et al. (2021) are the best and second performers, which consistently outperform the Top-k baseline by a considerable margin. Top-p mostly exhibits inferior recall comparing to the Top-k baseline, so does Eta-sampling at the risk level of 1. As for the stability represented by standard error of risks, Top-k sampling reaches the best scores in most cases. In comparison, Adaptive sampling and Mirostat deliver comparable standard error of risks to Top-k sampling, whereas Top-p sampling and Eta-sampling are again inferior. Considering both diversity and stability, Adaptive sampling and Mirostat are the top 2 adaptive methods to be recommended, and Top-p sampling shall be the last to be considered.
We also show in fig. 5 that larger models of the same family have higher average recall at the same risk level comparing to the smaller ones. This conforms to the fact that larger models better captures the human text distribution. Please note that our metrics doesn’t allow a direct comparison between different model families, mainly due to their different vocabulary sizes and tokenizers, e.g., Llama-3 has a 128,256 vocabulary size, while Llama-2 has only 32, 000 vocabulary size. Moreover, our metrics also explicitly exclude the source of risk within the optimal allowed set, which is heavily dependent on a LLM’s capacity.
| Question | gpt2xl |
| What happens to you if you eat watermelon seeds? | You will die. |
| Where did fortune cookies originate? | Fortune cookies originated in the United States. |
Although Top-p sampling is indeed also adaptive regarding the truncation position, we show that Top-p sampling have a inherent limitation. When a larger portion of the probability mass is concentrated in the first few tokens (this often indicates smaller entropy), a fixed cumulative probability threshold will cut a longer tail off, and vice versa. However, there’s merely a weak correlation between the entropy of the LLM’s prediction and optimal truncation values, see fig. 6.
| Methods | Mean(std) Accuracy | ||
| Avg. risk level 1 | Avg. risk level 5 | Avg. risk level 15 | |
| Greedy | 0.338 | ||
| Naïve | 0.421(0.004) | ||
| Top-k | 0.401(0.010) | 0.436(0.008) | 0.421(0.010) |
| Top-p | 0.355(0.013) | 0.378(0.011) | 0.389(0.012) |
| Adaptive | 0.395(0.012) | 0.424(0.011) | 0.421(0.009) |
| Eta | 0.388(0.005) | 0.401(0.013) | 0.413(0.026) |
| Mirostat | 0.413(0.010) | 0.425(0.013) | 0.425(0.009) |
5.4 Validation on TruthfulQA Benchmark
Although our evaluation protocol is grounded by the thorough design process with reasonable simplifications, we would like to verify its effectiveness in the real-world scenario using the TruthfulQA Benchmark Lin et al. (2021). We evaluate the performance of gpt2-xl model with each truncation sampling method at the average risk levels of 1, 5 and 15 respectively. The evaluation results are shown in table 2. For all the methods other than greedy decoding, we run 3 times at each average risk level and report the mean and standard deviation (parenthetical value).
It can be observed that greedy decoding falls far behind sampling-based decoding strategies, which conforms to the issue of likelihood-oriented decoding discussed in section 1, as well as the findings in recent studies Cobbe et al. (2021); Wang et al. (2023); Wang and Zhou (2024); Shi et al. (2024a). The examples in fig. 7 also explain the unsatisfactory performance of greedy decoding, i.e., the decoding paths of the corrected answers might be excluded after ignoring the non-peak likelihoods. Similarly, all the truncation sampling methods at the low risk level achieves lower accuracy comparing to Naive sampling, due to the over-truncation of the decoding paths. At the average risk level of 5, all the truncation sampling methods slightly improve their own accuracy. Top-k sampling, Adaptive sampling and Mirostat also reach comparable or slightly higher accuracy in comparison to Naive sampling. However, further increased average risk level (means improved average recall and thus diversity) doesn’t benefit the performance on TruthfulQA, which is plausible. Moreover, there exists a even stronger correlation between Risk SE (Standard Error of Risks) and TruthfulQA accuracy, validating the importance of stability when evaluating an adaptive decoding method. The strong correlation between TruthfulQA accuracy and our proposed average recall as well as standard error of risks at different risk levels validate the soundness and effectiveness of our evaluation method.
6 Conclusion
In this work, we propose a evaluation protocol to assess the intrinsic capacity of truncation sampling methods for open-ended text generation. Our evaluation enjoys the merit of being independent on parameter tuning for the curated tasks. Its effectiveness is further validated by the results on the open-ended text generation setup of TruthfulQA Benchmark. The evaluation results also serve as user reference for creative tasks.
7 Limitations
In this work, we focus on the truncation sampling methods specially designed for the open-ended text generation scenario. There exist many related decoding strategies, which aim at improving different aspects of LLMs. For example, a line of decoding strategies are proposed to alleviate Hallucination or improve the reasoning ability, e.g., Dola Chuang et al. (2023), Context-aware decoding Shi et al. (2024b), Contrastive decoding O’Brien and Lewis (2023) and etc. However, they are beyond the scope of this study and thus not included in the discussion. Although our study is only based on text data in English for clarity, the conclusion should be transferable to other languages as well.
8 Broader Impact
Our study on the intrinsic capacity of sampling methods and their appropriate parameters for open-text generation may further promote the application of LLMs in creative industries. There exists a potential risk that our provided findings might be abused for generating harmful or fake information. However, our study itself is neutral and the mentioned risk is a general issue that LLMs face. We call for the attention on AI-Safety in the community.
Appendix A Appendix
A.1 Complete Record of the Experiment Runs
The scores of the individual runs on TruthfulQA benchmark are recorded in table 3, and the means and standard errors of recalls and risks at all average risk levels are listed in table 4. Note that due to a fixed amount of computation budget, we search the corresponding parameter value for each truncation sampling method till the average risk is close enough to the predefined value, thus resulting in the variations of the average risks. However, such variations are negligible given the minor differences.
A.2 The Advantage of Probability-Independent Metrics
In this section, we explain the practical advantages of our proposed probability-independent recall and risk metrics. As can be seen in fig. 9, the empirical distribution aligns with the by gpt2-xl predicted distribution given the same prefix in general: most of the tokens which posses high likelihood in the prediction also has a high probability based on the word frequencies of our collected CP-Trie data. However, there exists two differences:
Some tokens with high likelihood according to gpt2-xl have much lower probability according to the empirical distribution. The ranking of each tokens w.r.t. probability also differ in the two distributions.
A few tokens which should be reasonable candidates (by manual check) have 0 probability according to the empirical distribution.
For the first issue, as discussed in section 3.2, there exists no ideal probabilities for each token, and the discrepancy is not solvable by simply increasing the size of the data. For example, the "perfect" probabilities of the candidate tokens "with" and "at" are undefined and could even be regarded as equivalently important for open-ended text generation.
The second difference highlights the reliability of LLMs, i.e., the tokens which are assigned high likelihoods are in most cases reasonable. Note that we ignore the risk within the estimated optimal allowed set by design: All the tokens are counted as reasonable till the last token which has non-zero empirical probability, when they are arranged in a descending order according to the predicted probabilities. Thus these tokens with zero probabilities in the empirical distribution will not affect our evaluation of risk, making our method robust to noises and insufficient data support.
| Methods | Evaluation Runs | Mean/Std | ||||||||||
| Run 1 at average risk levels | Run 2 at average risk levels | Run 3 at average risk levels | average risk levels | |||||||||
| 1 | 5 | 15 | 1 | 5 | 15 | 1 | 5 | 15 | 1 | 5 | 15 | |
| Greedy Decoding | 0.338 | |||||||||||
| Naive Sampling | 0.420 | 0.426 | 0.416 | 0.421(0.004) | ||||||||
| Top-k Sampling | 0.412 | 0.447 | 0.410 | 0.389 | 0.432 | 0.435 | 0.402 | 0.428 | 0.419 | 0.401(0.010) | 0.436(0.008) | 0.421(0.010) |
| Top-p Sampling | 0.337 | 0.370 | 0.382 | 0.367 | 0.393 | 0.379 | 0.362 | 0.370 | 0.405 | 0.355(0.013) | 0.378(0.011) | 0.389(0.012) |
| Adaptive Sampling | 0.403 | 0.416 | 0.433 | 0.403 | 0.416 | 0.419 | 0.378 | 0.440 | 0.411 | 0.395(0.012) | 0.424(0.011) | 0.421(0.009) |
| Eta Sampling | 0.395 | 0.419 | 0.442 | 0.387 | 0.394 | 0.419 | 0.382 | 0.389 | 0.379 | 0.388(0.005) | 0.401(0.013) | 0.413(0.026) |
| Mirostat | 0.424 | 0.417 | 0.430 | 0.399 | 0.443 | 0.433 | 0.415 | 0.414 | 0.412 | 0.413(0.010) | 0.425(0.013) | 0.425(0.009) |
| Method | GPT2-XL | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 15 | 1.029 (0.006) | 0.220(0.0006) | 64 | 5.040 (0.613) | 0.290 (0.017) | 184 | 14.983(1.781) | 0.340 (0.018) |
| Top-p | 0.5705 | 0.999 (0.015) | 0.170 (0.0005) | 0.746 | 5.011(2.129) | 0.240 (0.015) | 0.8555 | 15.022 (6.210) | 0.338 (0.016) |
| Adaptive | 9.5e-4 | 1.000 (0.006) | 0.252 (0.0007) | 0.00011 | 4.997 (0.679) | 0.339(0.018) | 2.5e-05 | 14.995 (2.241) | 0.413 (0.018) |
| Eta | 0.318 | 1.000 (0.013) | 0.198 (0.0005) | 0.011 | 4.945 (1.484) | 0.301 (0.016) | 0.001 | 14.998 (4.261) | 0.404 (0.017) |
| Mirostat | 4.425 | 0.999 (0.005) | 0.236 (0.0007) | 5.9475 | 5.001 (0.717) | 0.326 (0.018) | 6.76 | 14.982 (2.501) | 0.401 (0.018) |
| Method | Llama-2-7b | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 14 | 0.986 (0.126) | 0.226 (0.016) | 61 | 4.987 (0.587) | 0.296 (0.017) | 177 | 14.961 (1.722) | 0.369 (0.018) |
| Top-p | 0.54 | 0.999 (0.529) | 0.156 (0.012) | 0.7665 | 4.990 (2.331) | 0.254 (0.015) | 0.9 | 14.989 (6.208) | 0.400 (0.016) |
| Adaptive | 0.0011 | 1.051 (0.154) | 0.257 (0.016) | 0.00014 | 4.991 (0.856) | 0.364 (0.017) | 3.1e-5 | 14.995 (2.966) | 0.470 (0.017) |
| Eta | 0.512 | 1.000 (0.563) | 0.192 (0.014) | 0.023 | 5.007 (2.599) | 0.297 (0.016) | 0.002 | 13.487 (6.531) | 0.407 (0.017) |
| Mirostat | 4.253 | 1.000 (0.133) | 0.236 (0.016) | 5.82 | 4.993 (0.650) | 0.349 (0.018) | 6.628 | 15.022 (2.286) | 0.474 (0.017) |
| Method | Llama-3-8B | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 14 | 1.023 (0.128) | 0.228 (0.016) | 59 | 4.982 (0.576) | 0.290 (0.017) | 172 | 15.025 ( 1.701) | 0.346 ( 0.018) |
| Top-p | 0.5395 | 1.000 (0.451) | 0.154 (0.013) | 0.736 | 4.998 (2.061) | 0.224 (0.014) | 0.855 | 14.993 ( 5.770) | 0.326 ( 0.016) |
| Adaptive | 0.0011 | 1.133 (0.167) | 0.260 (0.017) | 0.00017 | 5.006 (0.787) | 0.343 (0.018) | 3.7e-5 | 15.007 ( 2.685) | 0.418 ( 0.018) |
| Eta | 0.673 | 1.000 (0.445) | 0.181 (0.014) | 0.029 | 5.009 (2.112) | 0.271 (0.016) | 0.002 | 15.012 ( 6.009) | 0.373 ( 0.017) |
| Mirostat | 4.24 | 1.001 (0.139) | 0.230 (0.016) | 5.8175 | 5.001 (0.804) | 0.318 (0.018) | 6.6925 | 14.996 ( 2.630) | 0.393 ( 0.018) |
| Method | Llama-3-70B | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 14 | 1.014 ( 0.127) | 0.230 ( 0.016) | 60 | 5.038 ( 0.581) | 0.295 ( 0.017) | 173 | 15.024 ( 1.695) | 0.352 ( 0.018) |
| Top-p | 0.5695 | 1.001 ( 0.502) | 0.158 ( 0.013) | 0.758 | 4.999 ( 2.386) | 0.237 ( 0.015) | 0.8705 | 14.960 ( 6.685) | 0.332 ( 0.016) |
| Adaptive | 0.0011 | 1.004 ( 0.137) | 0.263 ( 0.017) | 0.00014 | 5.013 ( 0.787) | 0.353 ( 0.018) | 3.16e-5 | 14.986 ( 2.778) | 0.424 ( 0.018) |
| Eta | 0.37 | 1.004 ( 0.137) | 0.263 ( 0.017 ) | 0.014 | 5.032 ( 2.231) | 0.295 ( 0.016) | 0.001 | 15.076 ( 6.265) | 0.398 ( 0.018) |
| Mirostat | 4.21 | 1.001 ( 0.138) | 0.230 ( 0.016 ) | 5.91 | 5.001 ( 0.708) | 0.332 ( 0.018 | 6.84 | 15.021 ( 2.193) | 0.417 ( 0.018) |
| Method | Llama-2-70b | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 14 | 1.002 ( 0.128) | 0.232 ( 0.016 ) | 60 | 4.982 ( 0.583) | 0.307 ( 0.017) | 174 | 14.964 ( 1.712) | 0.375 ( 0.018) |
| Top-p | 0.6535 | 0.999 ( 0.475) | 0.189 ( 0.013 ) | 0.8465 | 4.988 ( 2.136) | 0.316 ( 0.016) | 0.9395 | 15.019 ( 5.522) | 0.468 ( 0.016) |
| Adaptive | 0.0011 | 1.000 ( 0.142) | 0.269 ( 0.017 ) | 1.2e-4 | 4.995 ( 0.796) | 0.374 ( 0.017) | 2.3e-5 | 15.007 ( 2.697) | 0.485 ( 0.017) |
| Eta | 0.092 | 1.002 ( 0.304) | 0.236 ( 0.015 ) | 0.003 | 5.057 ( 1.590) | 0.378 ( 0.017) | 0.00021 | 15.001 ( 4.243) | 0.510 ( 0.017) |
| Mirostat | 4.16 | 1.001 ( 0.135) | 0.238 ( 0.016 | 5.7875 | 5.004 ( 0.684) | 0.353 ( 0.018) | 6.67 | 14.991 ( 2.125) | 0.478 ( 0.017) |
| Method | Mixtral-8x7B | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 15 | 1.028 ( 0.134) | 0.229 ( 0.016) | 63 | 4.978 ( 0.598) | 0.301 ( 0.017) | 183 | 14.967 ( 1.757) | 0.366 ( 0.018) |
| Top-p | 0.6505 | 1.000 ( 0.535) | 0.192 ( 0.014 ) | 0.8375 | 5.007 ( 2.423) | 0.303 ( 0.015) | 0.9325 | 14.966 ( 6.139) | 0.456 ( 0.016) |
| Adaptive | 0.00105 | 1.000 ( 0.148) | 0.265 ( 0.017 ) | 0.00011 | 4.994 ( 0.798) | 0.372 ( 0.018) | 2.1e-5 | 15.014 ( 2.802) | 0.476 ( 0.017) |
| Eta | 0.087 | 1.001 ( 0.335) | 0.241 ( 0.015 ) | 0.003 | 5.061 ( 1.822) | 0.375 ( 0.017) | 0.000215 | 14.991 ( 4.922) | 0.506 ( 0.017) |
| Mirostat | 4.2775 | 1.000 ( 0.143) | 0.238 ( 0.016) | 5.845 | 4.995 ( 0.710) | 0.346 ( 0.018) | 6.6875 | 14.998 ( 2.213) | 0.461 ( 0.018) |
| Method | Mistral-7B | ||||||||
| Parameter | Risk | Recall | Parameter | Risk | Recall | Parameter | Risk | Recall | |
| Top-k | 14 | 0.965 ( 0.126) | 0.224 ( 0.016) | 62 | 4.968 ( 0.596) | 0.297 ( 0.017) | 181 | 15.006 ( 1.759) | 0.364 ( 0.018) |
| Top-p | 0.6565 | 1.001 ( 0.539) | 0.194 ( 0.014) | 0.8375 | 4.996 ( 2.476) | 0.303 ( 0.016 ) | 0.9315 | 15.038 ( 6.315) | 0.447 ( 0.016) |
| Adaptive | 0.00105 | 1.001 ( 0.152) | 0.260 ( 0.016) | 0.000115 | 4.993 ( 0.809) | 0.364 ( 0.018) | 2.2e-5 | 14.999 ( 2.757) | 0.466 ( 0.017) |
| Eta | 0.075 | 0.997 ( 0.307) | 0.243 ( 0.015) | 0.003 | 4.640 ( 1.542) | 0.368 ( 0.017) | 0.000196 | 15.009 ( 4.712) | 0.505 ( 0.017) |
| Mirostat | 4.1825 | 1.000 ( 0.141) | 0.236 ( 0.016) | 5.8125 | 4.999 ( 0.721) | 0.345 ( 0.018) | 6.71 | 14.978 ( 2.213) | 0.468 ( 0.018) |

### v2

@@
 Abstract
-Sampling-based decoding strategies have been widely adopted for Large Language Models (LLMs) in numerous applications, which target a balance between diversity and quality via temperature tuning and tail truncation (e.g., top-k and top-p sampling).
-Considering the high dynamic range of the candidate next-tokens given different prefixes, recent studies propose to adaptively truncate the tail of LLM’s predicted distribution.
-Although improved results haven been reported with these methods on open-ended text generation tasks, the results are highly dependent on the curated truncation parameters and exemplar text.
-In this paper, we propose a systematic way to estimate the intrinsic capacity of a truncation sampling method by considering the trade-off between diversity and risk at each decoding step, based on our collected prefix tree which preserves the context of a full sentence.
-Our work provides a comprehensive comparison between existing truncation sampling methods, as well as their recommended parameters as a guideline for users.
+Sampling-based decoding strategies have been widely adopted for Large Language Models (LLMs) in numerous applications, targeting a balance between diversity and quality via temperature tuning and tail truncation.
+Considering the strong dependency of the candidate next tokens on different prefixes, recent studies propose to adaptively truncate the tail of LLMs’ predicted distribution.
+Although improved results have been reported with these methods on open-ended text generation tasks, the results are highly dependent on the curated parameters and the limited exemplar text.
+In this paper, we propose a systematic way to estimate the capacity of a truncation sampling method by considering the trade-off between diversity and risk at each decoding step, based on our collected prefix tree which preserves the context of a full sentence.
+Our work offers a comprehensive comparison of existing truncation sampling methods and serves as a practical user guideline for their parameter selection.
+Our code is available at github repository.
 Balancing Diversity and Risk in LLM Sampling: How to Select Your Method and Parameter for Open-Ended Text Generation
-Yuxuan Zhou1,3, Margret Keuper1,2, Mario Fritz3 1University of Mannheim, Germany 2Max Planck Institute for Informatics, Saarland Informatics Campus, Germany 3CISPA Helmholz Center for Information Security, Germany zhouyuxuanyx@gmail.com, keuper@mpi-inf.mpg.de, fritz@cispa.de
+Yuxuan Zhou1,2, Margret Keuper1, Mario Fritz2 1University of Mannheim, Germany 2CISPA Helmholtz Center for Information Security, Germany yuxuan.zhou@cispa.de, keuper@uni-mannheim.de, fritz@cispa.de
 1 Introduction
@@
 (2023); Team et al.
-(2023) have demonstrated incredible performance in a variety of applications, and the reliability of decoding strategies has become a critical concern, especially where diverse and coherent samples are desired.
+(2023) have demonstrated exceptional performance across a variety of applications, and the reliability of decoding strategies has become a critical concern.
 Previous works have revealed that likelihood-maximization such as beam search Fan et al.
@@
 (2020); Meister et al.
-(2022) produces degenerated text which contains repetitive loops and incoherent context, particularly in the open-ended tasks.
+(2022) produces degenerate text which contains repetitive loops and incoherent context, particularly in open-ended tasks.
 Therefore, sampling-based decoding strategies, e.g., Top-p Holtzman et al.
@@
 (2021) have been adopted to consider different aspects of the generated text.
-We reveal that there exist two underlying issues in the current evaluation setup, which might hinder the assessment of a sampling method’s practical significance in real-world applications:
-The improvement of one method against another may be simply due to a better tuned parameter for the targeted exemplar text: the performance of sampling methods is sensitive to their parameters, and parameter sweep is often operated on a extremely sparse grid due to the high computation cost.
+We reveal two underlying issues in the current evaluation, which hinder the assessment of a method’s significance in real-world applications:
+The improvement of one method over another might be simply due to a better tuned parameter for the targeted task: the performance of sampling methods is sensitive to their parameters, and parameter sweep is often operated on a extremely sparse grid due to the high computation cost.
 This is especially problematic considering the non-linear dependency between performance and parameters.
-Users are agnostic to the optimal parameters in real-world applications: Practically speaking, users often pick parameters based on their own need for the compromise between diversity and quality, after few number of tryouts.
-There exits no universal optimal hyperparamters in different scenarios and users are agnostic to the optimal hyerparameters for their own tasks.
-The above issues exactly indicate the need for an evaluation that allows for estimating the theoretical capacity of a sampling method, independent of hyperparamter tuning.
-Moreover, the second issue additionally highlights the need for identifying the sweet spots of existing sampling methods, which could serve as a general guideline for parameter selection when applying a method.
-Based on the above analysis, we propose a systematic way to assess the inherent adaptability of a sampling method in different contexts.
-First, we rearrange Wikipedia-English 111https://dumps.wikimedia.org/ data in the form of a word-level prefix tree structure, or the so-called Trie Fredkin (1960); Ghasemi et al.
+Users are agnostic to the optimal parameters in real-world applications: Practically speaking, users often pick parameters based on their own need for the compromise between diversity and quality, after a few tryouts.
+There exists no universal optimal paramters in different scenarios and users are agnostic to the optimal parameters for their own tasks.
+The above issues exactly indicate the need for an evaluation that allows for estimating the theoretical capacity of a truncation sampling method (how well it adapts to the variation in data supports given different prefixes), independent of hyperparamter tuning.
+Moreover, the second issue additionally highlights the need to identify the sweet spots of existing sampling methods, which could serve as a user guideline for practitioners.
+In light of the above analysis, we propose a systematic way to assess the inherent adaptability of a sampling method.
+First, we rearrange Wikipedia-English 111https://dumps.wikimedia.org/ data into a word-level prefix tree, known as a Trie Fredkin (1960); Ghasemi et al.
 (2019).
-As shown in fig.
-3, all possible words that appear after a given prefix in the dataset are collected together as the child nodes, and their preceding word is regarded as the parent node.
-Starting from "Begin of Sequence" and collecting the child nodes recursively, we are able to transform the full dataset into a single prefix tree.
-It is noteworthy that a n-gram Trie Jurafsky (2000) tends to produce overestimated data support size given a prefix Bengio et al.
-(2000), due to the lost of contextual information outside the truncation window, as shown in fig.
-1.
+It is noteworthy that a n-gram Trie Jurafsky (2000) tends to overestimate the data support size given a prefix Bengio et al.
+(2000), as shown in Figure 1.
 In a similar spirit to Ding et al.
-(2024), we intentionally construct the prefix tree with only sentence-starting n-grams to preserve the context of a full sentence.
-thus we refer to the collected data as Context-Preserving Trie (CP-Trie).
-Since the child nodes of the ancestral prefix define the data support, the recall of a sampling method w.r.t.
-data support is able to be computed.
-Though the training datasets of modern LLMs are significantly larger than Wikipedia dataset, such a data support serve as a reasonable lower-bound, especially when only sentence-level context are considered for evaluation, see section 4.1 for more details.
-Given the context-preserving prefix tree, we are able to evaluate the theoretical capacity of a sampling method, by examining the amount of tokens within and out of the data support with varying truncation parameter values.
-As can be seen in fig.
-2, the optimal truncation position vary drastically given different prefixes and Top-k sampling could be regarded as a baseline method with zero adaptability.
-Therefore, an effective adaptive truncation mechanism is supposed to better follow such a variation, so that improved diversity can be achieved without harming the quality.
-In summary, the contributions of our paper are as follows:
-We establish an intrinsic evaluation benchmark based on the Context-Preserving Prefix tree data, which allows for estimating the theoretical capacity of different sampling decoding methods via our proposed diversity and stability metrics.
-We conduct a comprehensive comparison of existing sampling approaches on our proposed benchmark, which serve as a guideline for choosing the truncation sampling methods and their corresponding parameter selection in real-world applications.
+(2024), we construct the prefix tree with only sentence-starting n-grams to preserve full sentence context, called Context-Preserving Trie (CP-Trie).
+Given the CP-Trie, we are able to estimate the theoretical capacity of a sampling method, by examining the amount of tokens within and out of the data support with varying truncation parameter values.
+As shown in Figure 2, the truncation positions, which exactly cover the full data supports, vary drastically given different prefixes and Top-k sampling could be regarded as a baseline method with zero adaptability.
+Therefore, an adaptive truncation method is supposed to better follow such a variation, so that improved diversity can be achieved without harming the quality.
+In summary, our contributions are as follows:
+We establish an intrinsic evaluation benchmark based on the collected CP-Trie, which allows for estimating the theoretical capacity of different sampling methods via thoroughly designed diversity and stability metrics.
+We conduct a comprehensive comparison of existing sampling approaches, which serves as a guideline for choosing a method and its parameter in real-world applications.
+We reveal that sampling-based decoding methods are underestimated in the existing study (Shi et al., 2024a) due to the difficulty in parameter selection, highlighting the merit of our evaluation protocol.
 2 Related Work
-In this section, we summarize the recently proposed sampling-based decoding strategies and the widely adopted benchmarks as well as metrics for open-ended text generation.
+In this section, we summarize recent sampling decoding strategies, along with common benchmarks and metrics for open-ended text generation.
 2.1 Sampling-based Decoding Methods
-Vanilla sampling suffers from the risk of obtaining incoherent tokens, thus truncation of the tail distribution has been heavily discussed to alleviate such an issue, e.g., Top-k (Radford et al.
-(2018); Fan et al.
-(2018)) and Top-p sampling (Holtzman et al.
-(2020)).
+Vanilla sampling suffers from the risk of obtaining incoherent tokens; thus, truncation of the tail distribution has been heavily discussed, e.g., Top-k (Radford et al., 2018; Fan et al., 2018) and Top-p sampling (Holtzman et al., 2020).
 However, a fixed k or p is problematic when considering the high dynamic range of next reasonable tokens, as pointed out in more recent studies on adaptive sampling methods: Mirostat Basu et al.
@@
 Locally Typical Sampling Meister et al.
-(2023) assumes the generated text to retain a similar entropy rate to that of human-generated text.
+(2023) assumes that the generated text should retain a similar entropy rate to that of human-generated text.
 Adaptive Decoding Zhu et al.
 (2024) proposes to keep the entropy of the truncated distribution close to the original entropy.
-Although these approaches have been empirically evaluated on open-ended text generation tasks with curated truncation parameters and exemplar text, there still lacks a comprehensive comparison of their adaptability in more general cases.
+Although these approaches have been demonstrated to be effective, their performance is highly dependent on the curated truncation parameters and the limited exemplar text.
 2.2 Evaluation of Sampling-based Decoding
-Benchmarks The commonly adopted benchmarks include story generation with WritingPrompts dataset Fan et al.
+Common benchmarks include story generation with WritingPrompts dataset Fan et al.
 (2018), document continuation with WikiText-103 dataset Merity et al.
@@
 (2021), Zipf’s coefficient Holtzman et al.
-(2020) (simple Unigram rank-frequency statistics) and Self-BLEU Zhu et al.
+(2020) (Unigram rank-frequency) and Self-BLEU Zhu et al.
 (2018).
@@
 (2018); Holtzman et al.
-(2020), lower perplexity of the generated text doesn’t necessarily indicate better quality.
+(2020), lower perplexity of the generated text does not necessarily indicate better quality.
 And Holtzman et al.
-(2020) suggested that the perplexity of the generated text is supposed to be close to that of the human text.
+(2020) suggested that the perplexity of the generated text should be close to that of the human text.
 MAUVE Pillutla et al.
 (2021) takes the trade-off between precision and recall into account, by comparing the learnt distribution from a text generation model to the distribution of human-written text using divergence frontiers.
-A recent study Shi et al.
-(2024a) provides a comprehensive evaluation on a large collection of tasks, which are mostly based on exemplar-based metrics.
+Shi et al.
+(2024a) provides a comprehensive evaluation on a large collection of tasks, mostly relying on exemplar-based metrics.
+However, we reveal that such evaluation is affected by the biases in the curated parameters and limited exemplar text, and our evaluation method is shown to alleviate such an issue.
 3 Revisiting Truncation Sampling
+We begin by revisiting the formulation of truncation sampling, followed by identifying the unresolved challenges in evaluating truncation sampling methods.
 3.1 Problem Formulation
 Definition 3.1.
-where denotes the allowed set comprising candidate next-tokens for a given prefix, and is the renormalization term.
-Given the Context-Preserving Trie of a reference dataset, we could compute the estimate of the optimal allowed set as follows :
+where denotes the allowed set of candidate next tokens at the position, given a sequence of tokens as prefix. is the renormalization term.
+Given the Context-Preserving Trie of a reference dataset, we can compute the estimate of the optimal allowed set as follows :
 Definition 3.2.
 Let be the allowed set after truncation given the prefix .
-The optimal allowed set corresponds to the allowed set with the minimum size, while covering the full data support.
+The approximated optimal allowed set corresponds to the allowed set with the minimum size, while covering the full data support for the token based on the Trie.
 It is the solution to the following objective function:
+Note that the above definition is designed to exclude the risk of obtaining OOD tokens before the cutoff Finlayson et al.
+(2024), because such type of risk is unsolvable by truncation and is rather determined by the capacity of the trained LLMs.
+However, such risk is less severe compared to that introduced by inappropriate truncation, since LLMs exhibit a significant capability in predicting the next token Touvron et al.
+(2023); Achiam et al.
+(2023); Jiang et al.
+(2023); Team et al.
+(2023) and most OOD samples reside in the tail distribution.
 3.2 Remaining Issues
 We reveal three major issues in the evaluation of truncation sampling.
-We first summarize the problem of directly using probability as quality metric, then show the choice of truncation parameter has a significant impact on the evaluation.
-Finally, we reveal that a minor difference of Recall and Risk values may result in significant changes in diversity and quality of the generated text.
+We first summarize the problem of directly using probability as quality metric, then show that the choice of truncation parameter has a significant impact on the evaluation.
 Unreliable Probability The probabilities of both the predicted and empirical distribution are not reliable for reflecting the quality of a text.
-Word frequencies are average statistics across various topics, and assuming the optimal probabilities or the optimal ranking of each reasonable next token is ill-posed.
-Impact of Parameter Selection We highlight the complexity and biases in parameter selection: Top-k and Top-p have constant upper bounds, i.e., the vocabulary size and 1, respectively.
+Word frequencies are average statistics across various topics, and the optimal probabilities or ranking of each next token is ill-posed.
+Parameter Sensitivity We highlight the complexity and biases in parameter selection: Top-k and Top-p have constant upper bounds, i.e., the vocabulary size and 1, respectively.
 In contrast, the upper bounds of -sampling and adaptive sampling are dependent on LLM’s predicted distribution, because they truncate the tail distribution based on the likelihood of tokens and the slope of Min-Max scaled entropy, respectively.
-The importance of identifying the effective ranges of such parameters is also reflected in the authors’ choice of numeral digit for their proposed parameters.
+The importance of identifying the effective ranges of such parameters is also reflected in the authors’ choice of numeral digit for their parameters.
 For example, is set to in Zhu et al.
@@
 In comparison, the adopted values for Top-p sampling are merely two digits after zero, such as .
-Our analysis also shows the significance of identifying the sweet spots of different sampling methods.
-The Butterfly Effect Although the top few samples possess the most probability mass, we reveal that a minor change in the size of the allowed set at each decoding step could lead to major differences in the quality of the generated text, due to LLMs’ auto-regressive nature.
-For example, the overall probability of at least one bad token to appear at the position in a sequence of length increases rapidly as increases (due to the product operator):
-Similarly, if only probability mass is assigned to additional tokens at each step, for a sequence of length , there will be chances of obtaining extra diverse samples.
+This shows the significance of identifying the sweet spots of different sampling methods.
 4 Method
-In this section, we derive our method for evaluating different sampling-based decoding strategies.
-To circumvent the reliability issue of the probabilities we merely check whether the predicted next-token is in or out of the data support.
+In this section, we derive our metrics for evaluating different sampling-based decoding strategies. The metrics are carefully designed to address the issues discussed in Section 3.2.
 4.1 Probability-Independent Metrics
-To quantify the diversity and quality of a sampling method based on CP-Trie, we define the Recall and Risk of a sampling method regarding a given prefix below:
+To circumvent the unreliable probability issue, we merely check whether the predicted next token is in or out of the data support.
+Specifically, we define Recall and Risk to quantify diversity and quality of a sampling method on a single node of CP-Trie:
 Definition 4.1.
 is dependent on the parameter selection for truncation, e.g., k value in Top-k sampling.
-When the allowed set is smaller than the optimal one after truncation, Recall is smaller than one and Risk is regarded as zero.
+When the allowed set is smaller than the approximated optimal allowed set after truncation, Recall is smaller than one and Risk is regarded as zero.
 With further increased size of the allowed set, Recall reaches one but Risk emerges.
-Since the sizes of reasonable sets vary drastically for different prefixes, it is not possible to always retain the optimal allowed set with a predefined parameter.
+Since the sizes of reasonable sets vary drastically for different prefixes, it is not possible to always retain the approximated optimal allowed set with a predefined parameter.
 In this case, we reveal that the adaptability w.r.t.
 the varying size of data support of a sampling method indeed determines its effectiveness in real-world application.
-Note that we ignore the risk of obtaining bad samples within the optimal allowed set, because such type of risk is unsolvable by truncation and is rather determined by the inherent capacity of the trained LLMs.
-However, such risk is less severe comparing to that introduced by inappropriate truncation, since LLMs exhibit a significant capability in predicting the next token Touvron et al.
-(2023); Achiam et al.
-(2023); Jiang et al.
-(2023); Team et al.
-(2023) and most bad samples reside in the tail distribution.
-More importantly, our evaluation doesn’t rely on the empirical probability, which is biased and inaccurate due to limited dataset size or context window size.
+More importantly, our evaluation does not rely on the empirical probability, which is biased and inaccurate due to limited dataset size or context window size.
 However, the tokens which appear in the dataset could be confidently regarded as reasonable, regardless of their actual probabilities.
-In addition, considering that temperature could change the flatness of distribution arbitrarily, we use ratio of token counts instead of probability mass to make the evaluation independent on temperature tuning and exemplar text.
-For a more detailed discussion with supporting examples, please refer to section A.2.
-4.2 Parameter-Independent Evaluation
-To eliminate the huge impact of parameter tuning on fair evaluation, we define the final diversity metric Recall at a given Risk level as follows:
+In addition, considering that temperature could change the flatness of distribution arbitrarily, we adopt ratio of token counts instead of probability mass to make the evaluation independent of temperature tuning and exemplar text.
+For a detailed discussion with supporting examples, please refer to Section A.2.
+4.2 Tuning-Independent Evaluation
+To eliminate the huge impact of Parameter Sensitivity issue on fair evaluation, we adopt Average Recall (AR) at an average Risk and Risk Standard Error (RSE) at an average Risk to quantify diversity and stability of a sampling method across nodes of CP-Trie, respectively:
 Definition 4.2.
-Analogously, a family of critical values such as can be easily defined.
-From the definition, it can be seen that such a metric is no longer dependent on the selection of parameter, thus it reflects the genuine capacity of a sampling method regardless of parameter tuning.
-This allows for a fair comparison between different sampling methods, especially considering their drastically different effective ranges, as mentioned in section 1 and section 3.2.
-4.3 The Priority of Low Variance in Risk
-Despite that more diverse text is desired for many generation tasks, the minimum requirement of the text to be coherent and reasonable is mostly prioritized in real-world applications.
-Besides the average recall at a given risk level, it is noteworthy that a stable adaptive truncation mechanism is also preferred , i.e., the variance of risks should be kept as low as possible at the given average risk level.
-Conjecture 4.3.
-At a given average risk level, the total amount of risk when generating a sequence of length is reduced with decreased variance of the risks at each decoding step.
-The above conjecture can be mainly understood by the auto-regressive generation process of LLMs, as shown in eq.
-3.
-The in-distribution probability dependent on the risk at each decoding step and is minimized when the product of the probabilities is maximized at a given average risk level.
-We could infer that the sum of the in-distribution probabilities is approximated determined at a given average risk level.
-For simplification, we assume their sum is unchanged.
-According to AM-GM inequality, the maximum of the product is achieved when each individual component is equal to each other.
-Roughly speaking, the less variance of the probability masses at each step, the larger their product and thus the smaller the total risk is.
+where the superscript denotes the node in the evaluation set of nodes on the prefix tree. Analogously, a family of critical values such as can be easily defined.
+Since is now determined by the given average Risk, the diversity metric reflects the genuine capacity of a sampling method regardless of parameter tuning.
+This allows for a fair comparison of different sampling methods, especially considering their drastically different effective ranges, as mentioned in Section 1 and Section 3.2.
 5 Experiment
-In this section, we conduct evaluation of existing sampling-based decoding approaches on our re-collected EnWiki CP-Trie dataset.
-We aim to estimate the inherent adaptability of exiting sampling-based methods and the results could be used as references for the application of LLMs in open-ended tasks.
+In this section, we conduct evaluation of existing sampling-based decoding approaches on our collected EnWiki CP-Trie dataset.
+We aim to estimate the inherent adaptability of sampling-based methods and the results could be used as references for the application of LLMs in open-ended tasks.
 5.1 Data Collection
-As mentioned above, the construction of the CP-Trie is a re-collection of an existing dataset.
-We apply the described procedure to the English subset of Wikipedia dataset and name the resulting dataset EnWiki CP-Trie.
-Although the core idea is straight-forward to understand, we elaborate the main design choices in the following:
-Basic Unit There are many possible units for splitting the datasets into individual fragments, such as article, paragraph, sentence and n-grams.
-Constructing a tree based on articles or paragraphs may require a larger amount of data than the training data of LLMs to guarantee an adequate number of branches (because LLMs lean to interpolate), whereas the construction based on n-grams suffers from poor contextual information and are heavily biased towards common tuplets of n tokens regardless of the context.
-Therefore, we adopt sentence as the basic unit for our dataset, which guarantees a coherent context at sentence-level and requires a much smaller amount of data than the training data.
-Filtering To avoid invalid words or rare proper names which are unreasonable for the model to predict, we exclude the sentences containing such words by checking their presence in the WORD LIST dataset, which is available on the website 222word-list dataset homepage.
-It contains a total amount of 354986 words and explicitly excludes proper names and compound words.
-Section titles are also excluded, because they are often incomplete sentences with poor contextual information.
-Statistics Wikipedia-English dataset has a total number of articles, which results in EnWiki CP-Trie with number of leaves after conversion, as shown in fig. 4.
-Storage The extracted prefix tree is implemented as a nested dictionary and saved in a single JSON file.
-Since each lookup at any depth has constant complexity, the retrieval from our dataset is highly efficient.
+We construct our Trie data based on the English subset of Wikipedia dataset, named EnWiki CP-Trie.
+As shown in Figure 3, all possible words that appear after a given prefix in the dataset are treated as child nodes, with their preceding word regarded as the parent node.
+Starting from "Begin of Sequence" and collecting the child nodes recursively, we are able to transform the full dataset into a single prefix tree.
+We elaborate the main design choices in the following:
+Basic Unit.
+It is possible to split the datasets into articles, paragraphs, sentences or n-grams.
+Constructing a tree based on articles or paragraphs may require more data than the training data of LLMs to guarantee an adequate number of branches (because LLMs lean to interpolate), whereas the construction based on n-grams suffers from poor contextual information and is heavily biased towards common tuplets of n tokens regardless of the context.
+Therefore, we adopt sentence as the basic unit, which guarantees a coherent context at sentence-level and requires much fewer data than training.
+It is noteworthy that a n-gram Trie Jurafsky (2000) tends to overestimate the data support size given a prefix Bengio et al.
+(2000), due to the loss of information outside the contextual window, as shown in Figure 1.
+Filtering.
+To avoid invalid words or rare proper names which are unreasonable for the model to predict, we exclude the sentences containing such words by checking their presence in the WORD LIST dataset, which is available on the website 222word-list dataset homepage.
+It contains 354986 words in total and explicitly excludes proper names and compound words.
+Section titles are also excluded, which are often incomplete sentences with poor contextual information.
+Statistics. Wikipedia-English dataset contains articles, which result in EnWiki CP-Trie with leaves, see Figure 4.
+Storage.
+The prefix tree is implemented as a nested dictionary and saved in JSON format.
+Since each lookup at any depth has constant complexity, the retrieval is highly efficient.
 Moreover, the dictionary is easily extendable if extra data are needed for a more accurate estimation of the full data support.
 5.2 Evaluation Setup
-Baselines Our evaluation includes Top-k sampling Radford et al.
+Baselines.
+Our evaluation includes Top-k sampling Radford et al.
 (2018); Fan et al.
@@
 (2021) into comparison.
-Evaluation Data To guarantee a tight lower bound of the ideal data support given different prefixes, we first sort the sub-nodes according to their total number of leaves at each depth, then we select the top 10 sub-trees with different sentence starting tokens for evaluation.
+Evaluation Data.
+To guarantee a tight lower bound of the ideal data support given different prefixes, we first sort the sub-nodes according to their total number of leaves at each depth, then we select the top 10 sub-trees with different sentence starting tokens for evaluation.
 Moreover, we keep the top 2 child nodes at each depth till depth 6, since the empirical data support becomes less adequate at large depth.
 This results in an evaluation set of prefixes with varying lengths in total.
-Evaluation Metrics As discussed in section 4.2, we measure the improvement in diversity via the increase of average recall at a given risk level, and the reduction of the total risk in the auto-regressive process via the decrease of standard deviation at a given risk level, also referred to as stability.
-LLMs To ensure that the obtained conclusion generalizes to different models, we adopt Llama-2 Touvron et al.
-(2023), Llama-3 Dubey et al.
-(2024), Mistral Jiang et al.
-(2023, 2024) families of different sizes and GPT-2-XL Radford et al.
+Evaluation Metrics.
+We measure the improvement in diversity via the increase of Average Recall(AR) at an average Risk, and the improvement of the stability at each decoding step in the auto-regressive process via the decrease of Risk Standard Error (RSE) at an average Risk.
+We adopt AR and RSE at average Risks of 1, 5 and 15 for comparison, representing low, medium, and high-risk regions, respectively.
+LLMs.
+To ensure that the conclusion generalizes to different models, we adopt Llama Touvron et al.
+(2023); Dubey et al.
+(2024) family, Mistral Jiang et al.
+(2023, 2024) family and GPT-2-XL Radford et al.
 (2019) for comparison.
-Implementation Our implementation mainly relies on Pytorch Paszke et al.
+Tokenization.
+Since different LLMs are trained with different encoding methods, the evaluation has to be independent of the encoding methods.
+We solve this issue by constructing the CP-Trie with either a word or punctuation.
+For example, if the predicted next token corresponds to “sec", which is a part of the in-distribution word "section", then we regard this as a correct prediction.
+The second part “tion" is regarded as a hidden child node and is skipped in the evaluation.
+Parameter Search.
+We apply grid search to determine the corresponding parameters of different sampling methods for each average Risk.
+To address the highly non-linear dependency between the sampling methods and their truncation parameters, we employ an efficient coarse-to-fine grid search strategy: the number of grids is initially set to 2000.
+If a parameter results in an average Risk within ±0.1 of the target value, it is considered a feasible solution.
+Otherwise, an additional grid search is performed within a smaller interval until a feasible solution is found, based on the initial search results.
+The grids are determined using Llama3-70B and are applied consistently across all models.
+As shown in Table 7, almost all the deviations in the average Risks are much smaller than 0.1, demonstrating the robustness of our strategy.
+Implementation.
+Our implementation mainly relies on Pytorch Paszke et al.
 (2017), HuggingFace Wolf et al.
 (2020) and OpenAI API 333https://pypi.org/project/openai/ library.
-We implement a truncation sampling method ourselves if the official implementation is not available.
-For all truncation methods, the minimum size of the allowed set is set to 1 to prevent breaking the sampling process.
-| Model | Method | Avg. risk level 1 | Avg. risk level 5 | Avg. risk level 15 | ||||||
-| Parameter | Risk Std Error | Recall | Parameter | Risk Std Error | Recall | Parameter | Risk Std Error | Recall | ||
-| GPT2-XL | Top-k | 15 | 0.006 | 0.220 | 64 | 0.613 | 0.290 | 184 | 1.781 | 0.340 |
+We implement a truncation sampling method ourselves if the official implementation is unavailable.
+For all methods, the minimum size of the allowed set is set to 1 to prevent breaking the sampling process.
+| Model | Method | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 | ||||||
+| Parameter | RSE | AR | Parameter | RSE | AR | Parameter | RSE | AR | ||
+| GPT2-XL | Adaptive | 9.5e-4 | 0.006 | 0.252 | 1.1e-4 | 0.679 | 0.339 | 2.5e-05 | 2.241 | 0.413 |
+| Mirostat | 4.425 | 0.005 | 0.236 | 5.9475 | 0.717 | 0.326 | 6.76 | 2.501 | 0.401 | |
+| Top-k | 15 | 0.006 | 0.220 | 64 | 0.613 | 0.290 | 184 | 1.781 | 0.340 | |
+| Eta | 0.318 | 0.013 | 0.198 | 0.011 | 1.484 | 0.301 | 0.001 | 4.261 | 0.404 | |
 | Top-p | 0.5705 | 0.015 | 0.170 | 0.746 | 2.129 | 0.240 | 0.8555 | 6.210 | 0.338 | |
-| Adaptive | 9.5e-4 | 0.006 | 0.252 | 1.1e-4 | 0.679 | 0.339 | 2.5e-05 | 2.241 | 0.413 | |
-| Eta | 0.318 | 0.013 | 0.198 | 0.011 | 1.484 | 0.301 | 0.001 | 4.261 | 0.404 | |
-| Mirostat | 4.425 | 0.005 | 0.236 | 5.9475 | 0.717 | 0.326 | 6.76 | 2.501 | 0.401 | |
-| Llama-2-7b | Top-k | 14 | 0.126 | 0.226 | 61 | 0.587 | 0.296 | 177 | 1.722 | 0.369 |
+| Llama-2-7b | Adaptive | 1.1e-3 | 0.154 | 0.257 | 1.4e-4 | 0.856 | 0.364 | 3.1e-5 | 2.966 | 0.470 |
+| Mirostat | 4.253 | 0.133 | 0.236 | 5.82 | 0.650 | 0.349 | 6.628 | 2.286 | 0.474 | |
+| Top-k | 14 | 0.126 | 0.226 | 61 | 0.587 | 0.296 | 177 | 1.722 | 0.369 | |
+| Eta | 0.512 | 0.563 | 0.192 | 0.023 | 2.599 | 0.297 | 0.002 | 6.531 | 0.407 | |
 | Top-p | 0.54 | 0.529 | 0.156 | 0.7665 | 2.331 | 0.254 | 0.9 | 6.208 | 0.400 | |
-| Adaptive | 1.1e-3 | 0.154 | 0.257 | 1.4e-4 | 0.856 | 0.364 | 3.1e-5 | 2.966 | 0.470 | |
-| Eta | 0.512 | 0.563 | 0.192 | 0.023 | 2.599 | 0.297 | 0.002 | 6.531 | 0.407 | |
-| Mirostat | 4.253 | 0.133 | 0.236 | 5.82 | 0.650 | 0.349 | 6.628 | 2.286 | 0.474 | |
-| Llama-2-70b | Top-k | 14 | 0.128 | 0.232 | 60 | 0.583 | 0.307 | 174 | 1.712 | 0.375 |
+| Llama-2-70b | Adaptive | 0.0011 | 0.142 | 0.269 | 1.2e-4 | 0.796 | 0.374 | 2.3e-5 | 2.697 | 0.485 |
+| Mirostat | 4.16 | 0.135 | 0.238 | 5.7875 | 0.684 | 0.353 | 6.67 | 2.125 | 0.478 | |
+| Top-k | 14 | 0.128 | 0.232 | 60 | 0.583 | 0.307 | 174 | 1.712 | 0.375 | |
+| Eta | 0.092 | 0.304 | 0.236 | 0.003 | 1.590 | 0.378 | 2.1e-4 | 4.243 | 0.510 | |
 | Top-p | 0.6535 | 0.475 | 0.189 | 0.8465 | 2.136 | 0.316 | 0.9395 | 5.522 | 0.468 | |
-| Adaptive | 0.0011 | 0.142 | 0.269 | 1.2e-4 | 0.796 | 0.374 | 2.3e-5 | 2.697 | 0.485 | |
-| Eta | 0.092 | 0.304 | 0.236 | 0.003 | 1.590 | 0.378 | 2.1e-4 | 4.243 | 0.510 | |
-| Mirostat | 4.16 | 0.135 | 0.238 | 5.7875 | 0.684 | 0.353 | 6.67 | 2.125 | 0.478 | |
-| Llama-3-8B | Top-k | 14 | 0.128 | 0.228 | 59 | 0.576 | 0.290 | 172 | 1.701 | 0.346 |
+| Llama-3-8B | Adaptive | 1.1e-3 | 0.167 | 0.260 | 1.7e-4 | 0.787 | 0.343 | 3.7e-5 | 2.685 | 0.418 |
+| Mirostat | 4.24 | 0.139 | 0.230 | 5.8175 | 0.804 | 0.318 | 6.693 | 2.630 | 0.393 | |
+| Top-k | 14 | 0.128 | 0.228 | 59 | 0.576 | 0.290 | 172 | 1.701 | 0.346 | |
+| Eta | 0.673 | 0.445 | 0.181 | 0.029 | 2.112 | 0.271 | 0.002 | 6.009 | 0.373 | |
 | Top-p | 0.5395 | 0.451 | 0.154 | 0.736 | 2.061 | 0.224 | 0.855 | 5.770 | 0.326 | |
-| Adaptive | 1.1e-3 | 0.167 | 0.260 | 1.7e-4 | 0.787 | 0.343 | 3.7e-5 | 2.685 | 0.418 | |
-| Eta | 0.673 | 0.445 | 0.181 | 0.029 | 2.112 | 0.271 | 0.002 | 6.009 | 0.373 | |
-| Mirostat | 4.24 | 0.139 | 0.230 | 5.8175 | 0.804 | 0.318 | 6.693 | 2.630 | 0.393 | |
-| Llama-3-70B | Top-k | 14 | 0.127 | 0.230 | 60 | 0.581 | 0.295 | 173 | 1.695 | 0.352 |
+| Llama-3-70B | Adaptive | 1.1e-3 | 0.137 | 0.263 | 1.4e-4 | 0.787 | 0.353 | 3.16e-5 | 2.778 | 0.424 |
+| Mirostat | 4.21 | 0.138 | 0.230 | 5.91 | 0.708 | 0.332 | 6.84 | 2.193 | 0.417 | |
+| Top-k | 14 | 0.127 | 0.230 | 60 | 0.581 | 0.295 | 173 | 1.695 | 0.352 | |
+| Eta | 0.37 | 0.137 | 0.263 | 0.014 | 2.231 | 0.295 | 0.001 | 6.265 | 0.398 | |
 | Top-p | 0.5695 | 0.502 | 0.158 | 0.758 | 2.386 | 0.237 | 0.8705 | 6.685 | 0.332 | |
-| Adaptive | 1.1e-3 | 0.137 | 0.263 | 1.4e-4 | 0.787 | 0.353 | 3.16e-5 | 2.778 | 0.424 | |
-| Eta | 0.37 | 0.137 | 0.263 | 0.014 | 2.231 | 0.295 | 0.001 | 6.265 | 0.398 | |
-| Mirostat | 4.21 | 0.138 | 0.230 | 5.91 | 0.708 | 0.332 | 6.84 | 2.193 | 0.417 | |
-| Mixtral-7B | Top-k | 14 | 0.126 | 0.224 | 62 | 0.596 | 0.297 | 181 | 1.759 | 0.364 |
+| Mixtral-7B | Adaptive | 0.00105 | 0.152 | 0.260 | 1.2e-4 | 0.809 | 0.364 | 2.2e-5 | 2.757 | 0.466 |
+| Mirostat | 4.1825 | 0.141 | 0.236 | 5.8125 | 0.721 | 0.345 | 6.71 | 2.213 | 0.468 | |
+| Top-k | 14 | 0.126 | 0.224 | 62 | 0.596 | 0.297 | 181 | 1.759 | 0.364 | |
+| Eta | 0.075 | 0.307 | 0.243 | 0.003 | 1.542 | 0.368 | 1.96e-4 | 4.712 | 0.505 | |
 | Top-p | 0.6565 | 0.539 | 0.194 | 0.8375 | 2.476 | 0.303 | 0.9315 | 6.315 | 0.447 | |
-| Adaptive | 0.00105 | 0.152 | 0.260 | 1.2e-4 | 0.809 | 0.364 | 2.2e-5 | 2.757 | 0.466 | |
-| Eta | 0.075 | 0.307 | 0.243 | 0.003 | 1.542 | 0.368 | 1.96e-4 | 4.712 | 0.505 | |
-| Mirostat | 4.1825 | 0.141 | 0.236 | 5.8125 | 0.721 | 0.345 | 6.71 | 2.213 | 0.468 | |
-| Mixtral-8x7B | Top-k | 15 | 0.134 | 0.229 | 63 | 0.598 | 0.301 | 183 | 1.757 | 0.366 |
+| Mixtral-8x7B | Adaptive | 0.00105 | 0.148 | 0.265 | 1.1e-4 | 0.798 | 0.372 | 2.1e-5 | 2.802 | 0.476 |
+| Mirostat | 4.2775 | 0.143 | 0.238 | 5.845 | 0.710 | 0.346 | 6.6875 | 2.213 | 0.461 | |
+| Top-k | 15 | 0.134 | 0.229 | 63 | 0.598 | 0.301 | 183 | 1.757 | 0.366 | |
+| Eta | 0.087 | 0.335 | 0.241 | 0.003 | 1.822 | 0.375 | 2.15e-4 | 4.922 | 0.506 | |
 | Top-p | 0.6505 | 0.535 | 0.192 | 0.8375 | 2.423 | 0.303 | 0.9325 | 6.139 | 0.456 | |
-| Adaptive | 0.00105 | 0.148 | 0.265 | 1.1e-4 | 0.798 | 0.372 | 2.1e-5 | 2.802 | 0.476 | |
-| Eta | 0.087 | 0.335 | 0.241 | 0.003 | 1.822 | 0.375 | 2.15e-4 | 4.922 | 0.506 | |
-| Mirostat | 4.2775 | 0.143 | 0.238 | 5.845 | 0.710 | 0.346 | 6.6875 | 2.213 | 0.461 | |
-5.3 Comparison at Different Risk Levels
-In this section, we conduct a comprehensive study of different truncation sampling methods at different risk levels.
+5.3 Comparison at Different Average Risks
+In this section, we conduct a comprehensive study of different truncation sampling methods at different average Risks.
 As discussed in Section 4.2, this allows for a fair comparison which is independent of parameter tuning.
-Moreover, we provide the corresponding parameters for each truncation sampling method at different risk levels, which could serve as user reference for the parameter selection of the compared methods.
-As can be seen in table 1, different truncation sampling methods are compared at the average risk level of 1, 5, and 15 respectively.
+Moreover, we provide the corresponding parameters for each truncation sampling method at different average Risks, which could serve as a user reference for these methods.
+As can be seen in Table 1, different truncation sampling methods are compared at the average Risk of 1, 5, and 15 respectively.
 As discussed in Section 4.1, our defined risk and recall metrics explicitly exclude the source of risk induced by a LLM’s capacity by design, thus similar parameter values correspond to the same risk level for most sampling methods across various model types and sizes.
-This exactly showcases the advantage of our evaluation being parameter-independent and sustainable to the rapid update of LLMs.
+This exactly showcases the advantage of our evaluation being tuning-independent and sustainable to the rapid update of LLMs.
 Among the evaluated methods, Eta-sampling Hewitt et al.
-(2022) is the most sensitive to the changes of model type and size especially at risk levels of 1 and 5, which might hinder its practical significance at a low risk level.
-Regarding diversity, i.e., the average recall at the same average risk level, Adaptive sampling Zhu et al.
+(2022) is sensitive to the changes of model type and size, which might hinder its practical significance especially at a low risk level.
+Regarding diversity, i.e., the average recall at the same average Risk, Adaptive sampling Zhu et al.
 (2024) and Mirostat Basu et al.
 (2021) are the best and second performers, which consistently outperform the Top-k baseline by a considerable margin.
-Top-p mostly exhibits inferior recall comparing to the Top-k baseline, so does Eta-sampling at the risk level of 1.
+Top-p mostly exhibits inferior recall comparing to the Top-k baseline, so does Eta-sampling at the average Risk of 1.
 As for the stability represented by standard error of Risks, Top-k sampling reaches the best scores in most cases.
 In comparison, Adaptive sampling and Mirostat deliver comparable standard error of risks to Top-k sampling, whereas Top-p sampling and Eta-sampling are again inferior.
-Considering both diversity and stability, Adaptive sampling and Mirostat are the top 2 adaptive methods to be recommended, and Top-p sampling shall be the last to be considered.
-We also show in fig.
-5 that larger models of the same family have higher average recall at the same risk level comparing to the smaller ones.
+Considering both diversity and stability, Adaptive sampling and Mirostat are the top 2 adaptive methods to be recommended, whereas Top-p sampling shall be the last two methods to be considered.
+We also show in Figure 5 that larger models of the same family have higher average recall at the same risk level comparing to the smaller ones.
 This conforms to the fact that larger models better captures the human text distribution.
-Please note that our metrics doesn’t allow a direct comparison between different model families, mainly due to their different vocabulary sizes and tokenizers, e.g., Llama-3 has a 128,256 vocabulary size, while Llama-2 has only 32, 000 vocabulary size.
+Please note that our metrics does not allow a direct comparison between different model families, mainly due to their different vocabulary sizes and tokenizers, e.g., Llama-3 has a 128,256 vocabulary size, while Llama-2 has only 32, 000 vocabulary size.
 Moreover, our metrics also explicitly exclude the source of risk within the optimal allowed set, which is heavily dependent on a LLM’s capacity.
-| Question | gpt2xl |
-| What happens to you if you eat watermelon seeds? | You will die. |
-| Where did fortune cookies originate? | Fortune cookies originated in the United States. |
-Although Top-p sampling is indeed also adaptive regarding the truncation position, we show that Top-p sampling have a inherent limitation.
-When a larger portion of the probability mass is concentrated in the first few tokens (this often indicates smaller entropy), a fixed cumulative probability threshold will cut a longer tail off, and vice versa.
-However, there’s merely a weak correlation between the entropy of the LLM’s prediction and optimal truncation values, see fig.
-6.
 | Methods | Mean(std) Accuracy | ||
-| Avg. risk level 1 | Avg. risk level 5 | Avg. risk level 15 | |
+| Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 | |
 | Greedy | 0.338 | ||
@@
 (2021).
-We evaluate the performance of gpt2-xl model with each truncation sampling method at the average risk levels of 1, 5 and 15 respectively.
-The evaluation results are shown in table 2.
+The evaluation results using gpt2-xl are shown in Table 2.
 For all the methods other than greedy decoding, we run 3 times at each average risk level and report the mean and standard deviation (parenthetical value).
@@
 (2024a).
-The examples in fig.
-7 also explain the unsatisfactory performance of greedy decoding, i.e., the decoding paths of the corrected answers might be excluded after ignoring the non-peak likelihoods.
-Similarly, all the truncation sampling methods at the low risk level achieves lower accuracy comparing to Naive sampling, due to the over-truncation of the decoding paths.
+All the truncation sampling methods at the low risk level achieves lower accuracy comparing to Naive sampling, due to the over-truncation of the decoding paths.
 At the average risk level of 5, all the truncation sampling methods slightly improve their own accuracy.
 Top-k sampling, Adaptive sampling and Mirostat also reach comparable or slightly higher accuracy in comparison to Naive sampling.
-However, further increased average risk level (means improved average recall and thus diversity) doesn’t benefit the performance on TruthfulQA, which is plausible.
+However, further increased average risk level (means improved average recall and thus diversity) does not benefit the performance on TruthfulQA, which is plausible.
 Moreover, there exists a even stronger correlation between Risk SE (Standard Error of Risks) and TruthfulQA accuracy, validating the importance of stability when evaluating an adaptive decoding method.
-The strong correlation between TruthfulQA accuracy and our proposed average recall as well as standard error of risks at different risk levels validate the soundness and effectiveness of our evaluation method.
+The strong correlation between TruthfulQA accuracy and our proposed average recall as well as standard error of risks at different average Risks validate the soundness and effectiveness of our evaluation method.
+6 Revisiting Existing Evaluation
+In this section, we revisit the recent study (Shi et al., 2024a) by comparing sampling-based decoding methods at the same average Risks.
+We adopt the official implementation of Shi et al.
+(2024a).
+Following their setups, we adopt Llama-2-7B on MBPP (Austin et al., 2021), HumanEval (Austin et al., 2021) and GSM8K (Cobbe et al., 2021) to evaluate coding and math problem solving performance.
+Mean and standard deviation for three runs are reported in Table 3, Table 4 and Table 5, respectively.
+For all the three tasks, Mirostat does not perform well in general, probably because it is based on the Zipf-law of natural language and thus not suitable for code and math tasks.
+Notably, our greedy decoding baseline achieves significantly lower result than reported by Shi et al.
+(2024a) on HumanEval.
+Our results should be plausible, because the instruction tuned Llama-2-7B only achieves 7.9 according to Meta-Llama Github444 https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md .
+While their study concludes that deterministic methods outperform sampling methods across most tasks, our evaluation reveals that sampling methods are indeed underestimated.
+In contrast to the conclusion in Shi et al.
+(2024a), all the sampling-based decoding methods could achieve better performance than greedy decoding on HumanEval in Table 4.
+In addition, Top-p and eta sampling also beat greedy decoding at a low average Risk on GSM8K in Table 5.
+This observation underscores the challenges in parameter selection for sampling-based decoding, which is effectively addressed by our method.
+| Methods | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 |
+| Top-k | 19.70 (0.50) | 21.00 (2.30) | 20.50 (0.30) |
+| Top-p | 21.50 (1.30) | 21.10 (0.40) | 21.70 (0.70) |
+| Mirostat | 9.50 (0.30) | 8.80 (2.00) | 8.80 (0.40) |
+| Eta | 22.10 (0.70) | 19.10 (0.40) | 19.70 (0.40) |
+| Greedy | 24.00 | ||
+| Methods | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 |
+| Top-k | 5.68 (2.00) | 5.08 (1.52) | 6.50 (0.76) |
+| Top-p | 3.46 (1.05) | 5.89 (0.76) | 6.52 (2.43) |
+| Mirostat | 3.25 (0.76) | 4.27 (1.00) | 4.68 (0.58) |
+| Eta | 2.64 (2.01) | 6.91 (1.04) | 6.10 (1.32) |
+| Greedy | 2.44 | ||
+| Methods | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 |
+| Top-k | 7.56 (5.39) | 11.90 (0.80) | 11.73 (0.57) |
+| Top-p | 14.13 (0.47) | 8.72 (6.18) | 11.67 (0.11) |
+| Mirostat | 5.46 (0.47) | 5.74 (0.64) | 3.46 (2.10) |
+| Eta | 13.72 (0.46) | 8.42 (5.54) | 11.22 (0.75) |
+| Greedy | 13.19 | ||
 7 Conclusion
-In this work, we propose a evaluation protocol to assess the intrinsic capacity of truncation sampling methods for open-ended text generation.
-Our evaluation enjoys the merit of being independent on parameter tuning for the curated tasks.
-Its effectiveness is further validated by the results on the open-ended text generation setup of TruthfulQA Benchmark.
-The evaluation results also serve as user reference for creative tasks.
+In this work, we propose an evaluation protocol to assess the trade-off between diversity and quality of truncation sampling methods for open-ended text generation.
+Our evaluation enjoys the merit of being independent of parameter tuning for the curated tasks.
+The evaluation results also serve as a user reference for different downstream tasks.
 8 Limitations
@@
 There exist many related decoding strategies, which aim at improving different aspects of LLMs.
-For example, a line of decoding strategies are proposed to alleviate Hallucination or improve the reasoning ability, e.g., Dola Chuang et al.
+For example, a line of decoding strategies is proposed to alleviate hallucination or improve the reasoning ability, e.g., Dola Chuang et al.
 (2023), Context-aware decoding Shi et al.
 (2024b), Contrastive decoding O’Brien and Lewis (2023) and etc.
-However, they are beyond the scope of this study and thus not included in the discussion.
-Although our study is only based on text data in English for clarity, the conclusion should be transferable to other languages as well.
+However, these methods are beyond the scope of sampling-based decoding in this study and thus not included in the discussion.
+Although our study is only based on text data in English for clarity, the dataset can be extended to include other languages in the future.
+Due to time and resource constraints, we did not include all existing sampling-based decoding methods, such as Locally Typical Sampling Meister et al.
+(2023) and Min-P Sampling Nguyen et al.
+(2024), in our comparison.
+However, our benchmark is publicly available, and we plan to continuously update it with evaluations of additional methods in the future.
 9 Broader Impact
-Our study on the intrinsic capacity of sampling methods and their appropriate parameters for open-text generation may further promote the application of LLMs in creative industries.
+Our study on the capacity of sampling methods and their appropriate parameters for open-ended text generation may further promote the application of LLMs in creative industries.
 There exists a potential risk that our provided findings might be abused for generating harmful or fake information.
@@
 A.1 Complete Record of the Experiment Runs
-The scores of the individual runs on TruthfulQA benchmark are recorded in table 3, and the means and standard errors of recalls and risks at all average risk levels are listed in table 4.
+| Methods | Evaluation Runs | Mean/Std | ||||||||||
+| Run 1 at average Risks | Run 2 at average Risks | Run 3 at average Risks | average Risks | |||||||||
+| 1 | 5 | 15 | 1 | 5 | 15 | 1 | 5 | 15 | 1 | 5 | 15 | |
+| Greedy Decoding | 0.338 | |||||||||||
+| Naive Sampling | 0.420 | 0.426 | 0.416 | 0.421(0.004) | ||||||||
+| Top-k Sampling | 0.412 | 0.447 | 0.410 | 0.389 | 0.432 | 0.435 | 0.402 | 0.428 | 0.419 | 0.401(0.010) | 0.436(0.008) | 0.421(0.010) |
+| Top-p Sampling | 0.337 | 0.370 | 0.382 | 0.367 | 0.393 | 0.379 | 0.362 | 0.370 | 0.405 | 0.355(0.013) | 0.378(0.011) | 0.389(0.012) |
+| Adaptive Sampling | 0.403 | 0.416 | 0.433 | 0.403 | 0.416 | 0.419 | 0.378 | 0.440 | 0.411 | 0.395(0.012) | 0.424(0.011) | 0.421(0.009) |
+| Eta Sampling | 0.395 | 0.419 | 0.442 | 0.387 | 0.394 | 0.419 | 0.382 | 0.389 | 0.379 | 0.388(0.005) | 0.401(0.013) | 0.413(0.026) |
+| Mirostat | 0.424 | 0.417 | 0.430 | 0.399 | 0.443 | 0.433 | 0.415 | 0.414 | 0.412 | 0.413(0.010) | 0.425(0.013) | 0.425(0.009) |
+The scores of the individual runs on TruthfulQA benchmark are recorded in Table 6, and the means and standard errors of recalls and risks at all average Risks are listed in Table 7.
 Note that due to a fixed amount of computation budget, we search the corresponding parameter value for each truncation sampling method till the average risk is close enough to the predefined value, thus resulting in the variations of the average risks.
 However, such variations are negligible given the minor differences.
+Although Top-p sampling is indeed also adaptive regarding the truncation position, we show that Top-p sampling have a inherent limitation.
+When a larger portion of the probability mass is concentrated in the first few tokens (this often indicates smaller entropy), a fixed cumulative probability threshold will cut a longer tail off, and vice versa.
+However, there’s merely a weak correlation between the entropy of the LLM’s prediction and optimal truncation values, see Figure 7.
 A.2 The Advantage of Probability-Independent Metrics
 In this section, we explain the practical advantages of our proposed probability-independent recall and risk metrics.
-As can be seen in fig.
-9, the empirical distribution aligns with the by gpt2-xl predicted distribution given the same prefix in general: most of the tokens which posses high likelihood in the prediction also has a high probability based on the word frequencies of our collected CP-Trie data.
+As can be seen in Figure 8, the empirical distribution aligns with the by gpt2-xl predicted distribution given the same prefix in general: most of the tokens which posses high likelihood in the prediction also has a high probability based on the word frequencies of our collected CP-Trie data.
 However, there exists two differences:
@@
 Thus these tokens with zero probabilities in the empirical distribution will not affect our evaluation of risk, making our method robust to noises and insufficient data support.
-| Methods | Evaluation Runs | Mean/Std | ||||||||||
-| Run 1 at average risk levels | Run 2 at average risk levels | Run 3 at average risk levels | average risk levels | |||||||||
-| 1 | 5 | 15 | 1 | 5 | 15 | 1 | 5 | 15 | 1 | 5 | 15 | |
-| Greedy Decoding | 0.338 | |||||||||||
-| Naive Sampling | 0.420 | 0.426 | 0.416 | 0.421(0.004) | ||||||||
-| Top-k Sampling | 0.412 | 0.447 | 0.410 | 0.389 | 0.432 | 0.435 | 0.402 | 0.428 | 0.419 | 0.401(0.010) | 0.436(0.008) | 0.421(0.010) |
-| Top-p Sampling | 0.337 | 0.370 | 0.382 | 0.367 | 0.393 | 0.379 | 0.362 | 0.370 | 0.405 | 0.355(0.013) | 0.378(0.011) | 0.389(0.012) |
-| Adaptive Sampling | 0.403 | 0.416 | 0.433 | 0.403 | 0.416 | 0.419 | 0.378 | 0.440 | 0.411 | 0.395(0.012) | 0.424(0.011) | 0.421(0.009) |
-| Eta Sampling | 0.395 | 0.419 | 0.442 | 0.387 | 0.394 | 0.419 | 0.382 | 0.389 | 0.379 | 0.388(0.005) | 0.401(0.013) | 0.413(0.026) |
-| Mirostat | 0.424 | 0.417 | 0.430 | 0.399 | 0.443 | 0.433 | 0.415 | 0.414 | 0.412 | 0.413(0.010) | 0.425(0.013) | 0.425(0.009) |
 | Method | GPT2-XL | ||||||||